MOLECULAR OUTFLOWS FROM MASSIVE PROTOSTARS

\underline{KEE-TAE KIM}, Department of Astronomy, The University of Illinois, 1002 W. Grenn St., Urbana, IL 61801; E. CHURCHWELL, Department of Astronomy, The University of Wisconsin at Madison, 475 North Charter St., Madison, WI 53705; D. N. FRIEDEL, Department of Astronomy, The University of Illinois, 1002 W. Grenn St., Urbana, IL 61801; and M. SEWILO, Department of Astronomy, The University of Wisconsin at Madison, 475 North Charter St., Madison, WI 53705.

We carried out high-resolution (∼3′′) observations of 4 massive protostars in a bunch of molecular lines and 3 mm continuum using the BIMA array. Single compact (∼2′), poorly collimated bipolar molecular outflows are seen in lowresolution (∼30′′) single-dish maps of the 4 regions. But our BIMA array maps show multiple outflows in each region. The individual outflows have distinctly separated blue- and red-lobes and are better collimated than those seen in the single-dish maps. This result suggests that the poor collimation of massive molecular outflows shown by most previous single-dish studies may be caused by a combination of multiple outflows and/or low resolution. We will discuss the implications for massive star formation.